# Distilling the Essence: Efficient Reasoning Distillation via Sequence Truncation

### v1

@@ A.3 Loss Analysis
A.3.1 Openthoughts and Bespoke
Figures 13 and 14 present the loss curves comparing Left 50% (i.e., LSP 50%), Right 50%, and LSP 100% across all student-teacher pairs, while Tables 7 and 8 quantify the relative differences.
LSP 50% vs. LSP 100% Training on the first 50% of tokens consistently yields higher train loss than training on full sequences, for both soft and hard objectives. This arises because autoregressive language models predict each token conditioned on all preceding tokens; later positions benefit from richer context and are thus easier to predict. Since the reported loss $\mathcal{L}$ is averaged over all unmasked token positions $\mathcal{T}$,
@@ A.3.1 Openthoughts and Bespoke
However, we observe that the loss values between LSP 50% and LSP 100% are fairly close. The relative difference is smaller than the gap between the Left 50% and Right 50%, and much smaller than the difference between Right 50% and LSP 100%. This proximity indicates that early tokens provide a training signal that is nearly as effective as the full sequence. Given that the Left 50% subset concentrates on the more challenging reasoning steps (lacking many low-entropy ‘easy’ tokens that would typically lower the average loss), the fact that its loss remains competitive suggests that the early, high-entropy tokens are sufficient to drive significant alignment with both the teacher’s reasoning and the ground-truth distribution. Consequently, training on the first 50% captures the bulk of the learning dynamics, implying that the additional computational overhead of processing the latter half yields diminishing returns.
Left 50% vs. Right 50% Training on the first half of each sequence yields markedly lower soft and hard losses than training on the second half. The Left 50% configuration achieves substantially lower loss than the Right 50%, with relative differences consistently negative across different models and loss types. This significant gap can be attributed to (1) early tokens provide stronger learning signal and (2) that the right 50% starts with tokens midway through a reasoning sequence. This abrupt start deprives the model of the necessary context established in the first half, making subsequent token predictions harder and leading to higher losses.
@@
4B-32B Soft -9.20 -6.76 +9.04 +11.65
4B-32B Hard -15.44 -13.06 +8.18 +8.35
4B-8B Soft -14.93 -14.85 +8.99 +10.27
4B-8B Hard -15.68 -13.40 +8.66 +8.76
8B-32B Soft -9.75 -8.14 +9.09 +11.32
8B-32B Hard -16.89 -15.41 +7.43 +7.03
@@
4B-32B Soft -11.73 -14.34 +7.69 +5.86
4B-32B Hard -19.17 -21.49 +6.14 +2.51
4B-8B Soft -15.45 -18.05 +7.97 +7.08
4B-8B Hard -18.76 -21.66 +6.98 +3.63
8B-32B Soft -12.28 -14.93 +7.85 +6.29
8B-32B Hard -20.39 -22.48 +5.61 +2.55
@@
A.3.2 SkyT1
We omit the budget location ablation (Left vs. Right) for SkyT1 and focus on the comparison between partial (LSP 50%) and full-sequence (LSP 100%) training.
LSP 50% vs. LSP 100% As shown in Figure 15 and Table 9, the loss trends for SkyT1 mirror those of OpenThoughts and Bespoke. Training on the initial segment (LSP50%) results in consistently higher loss compared to the full sequence, reinforcing that earlier tokens are harder to predict. Despite this higher loss, the competitive performance of the LSP 50% model suggests that the early, high-entropy tokens are sufficient to drive significant alignment, capturing the bulk of the learning dynamics with less computational overhead.
Model Loss Type Avg (%) Last (%)
4B-32B Soft Loss +9.11 +7.68
4B-32B Hard Loss +6.40 +1.95
4B-8B Soft Loss +8.96 +6.66
4B-8B Hard Loss +6.90 +2.14
8B-32B Soft Loss +9.25 +7.60
8B-32B Hard Loss +5.48 +1.03
@@ A.3.2 SkyT1
A.4 Experimental Results of SkyT1
In this section, we present the experimental results for the SkyT1 dataset. We analyze SkyT1 separately from OpenThoughts and Bespoke due to its different linguistic characteristics. As shown in Table 10, models trained with LSP 50% retain, on average, approximately 85% of the performance achieved by full-sequence training. This substantial retention reinforces our finding that early tokens carry the strongest learning signal, enabling effective distillation even with truncated sequences. However, this retention rate is slightly lower than that observed for OpenThoughts and Bespoke.
LSP 50% LSP 100% Perf. Retention (%)
Pair AIME24 AIME25 AIME24 AIME25 AIME24 AIME25
32B-4B 0.1771 0.1844 0.2026 0.2057 87.4 89.6
8B-4B 0.1755 0.1896 0.2333 0.2130 75.2 89.0
32B-8B 0.2224 0.2135 0.2776 0.2583 86.1 82.7
Table 10: Performance comparison between LSP 50% and LSP 100% on AIME24 and AIME25 benchmarks for SkyT1. “Perf. Retention (%)” shows the performance retention of LSP 50% compared to LSP 100%
Two main factors likely contribute to this difference, as detailed in Appendix A.2.2. First, the first correct answer derivation occurs later in SkyT1. It typically occurs at approximately 54% of the full sequence length, which is significantly later than the $\approx 42-43\%$ observed for the Openthoughts and Bespoke datasets. Second, SkyT1 distributes reasoning more broadly across the sequence. Unlike OpenThoughts and Bespoke, where the answer section is typically a redundant summary of the Chain-of-Thought (CoT), the SkyT1 answer section sometimes continues or corrects the reasoning process in CoT section. Our linguistic analysis reveals that the answer section is fully entailed by the CoT in only 73% of SkyT1 examples, compared to 89% and 93% for OpenThoughts and Bespoke, respectively. We observed instances where the CoT derives an incorrect result (e.g. 47) that is subsequently corrected in the answer section (e.g. changed to 54). Because the answer section resides at the end of the sequence, it is excluded by the LSP 50% budget. Consequently, the model misses these critical corrections, leading to the slightly lower performance retention.
A.5 Ablation Studies
Our ablation objective in this section is to demonstrate that the model performance stems from the method itself rather than specific hyperparameter tuning. While we adopt $\lambda=0.5$ as the main experimental configuration, Table 11 shows that performance remains consistent across the range $\lambda\in\{0.25,0.5,0.75\}$ for 8B-4B teacher-student pair. The similarity in results among these $\lambda$ values indicates that the choice of 0.5 is not the sole driver of success, but rather that the method performs reliably provided that a balance between hard and soft losses is maintained.
This stability across intermediate $\lambda$ values is accompanied by a clear performance advantage over the pure supervised fine-tuning (SFT) setting ($\lambda=0$). All tested combined configurations ($\lambda\in\{0.25,0.5,0.75\}$) yield higher accuracy than the setting where soft loss is ignored, except for one setting when $\lambda=0.75$ where it is slightly lower than $\lambda=0$ in AIME25 but remains fairly close. Specifically, using our chosen $\lambda=0.5$ results in significant relative improvements: approximately 26.8% (AIME24) and 19.5% (AIME25) for LSP=0.5, and 8.9% (AIME24) and 9.7% (AIME25) for LSP=1.0.
@@ A.5 Ablation Studies
A.6 Miscellaneous
A.6.1 Compute Resources
@@
The second research question investigates compute-efficient training via sequence truncation and budget allocation. This involves two sets of experiments. First, we analyze the Sequence-Length Scaling Behavior by varying the Lead-Span Proportion (LSP) across $10$ levels: $p\in\{0.1,0.2,\dots,1.0\}$, representing the fraction of tokens retained from the start of the sequence. This is conducted on all $7$ teacher-student pairs for OpenThoughts and the $3$ largest pairs for Bespoke. Second, we perform a Budget Location Ablation to compare the Left 50% versus Right 50% of tokens. Since the Left 50% configuration is identical to LSP 0.5 (trained and evaluated in the first set), we reuse those results and only perform additional training for the Right 50% configuration. The total runs ($R_{RQ2}$) are aggregated from the Sequence-Length Scaling ($R_{RQ2}^{\text{Scale}}$) and Budget Location Ablation ($R_{RQ2}^{\text{Loc}}$) components as follows:
@@
RQ2 106 15,264 7,208 22,472
Total 142 20,448 9,656 30,104
A.6.2 Distillation Method Choice

### v2

@@ A.3 Loss Analysis
Figures 17181920, and 21 present the loss curves comparing partial-sequence training (LSP 50%) and full-sequence training (LSP 100%) across all student–teacher pairs for the five datasets (OpenThoughts, Bespoke, Synthetic-1, Nemotron, and SkyT1), respectively. For OpenThoughts and Bespoke, we additionally include the budget-location ablation comparing Left 50% (i.e., LSP 50%) and Right 50%. Tables 78910, and 11 quantify the relative differences.
LSP 50% vs. LSP 100% Across all five datasets, training on the first 50% of tokens consistently yields higher train loss than training on full sequences, for both soft and hard objectives. This arises because autoregressive language models predict each token conditioned on all preceding tokens; later positions benefit from richer context and are thus easier to predict. Since the reported loss $\mathcal{L}$ is averaged over all unmasked token positions $\mathcal{T}$,
@@
However, we observe that the loss values between LSP 50% and LSP 100% are fairly close across all five datasets: the relative differences remain modest throughout training, indicating that the partial-sequence objective tracks the full-sequence objective closely. Moreover, for datasets where we can compare Left vs. Right budget location, the Left 50% vs. LSP 100% gap is much smaller than the Right 50% vs. LSP 100% gap. This proximity suggests that the early tokens provide a strong and informative training signal. Given that the Left 50% subset concentrates on the more challenging reasoning steps (lacking many low-entropy ‘easy’ tokens that would typically lower the average loss), the fact that its loss remains competitive suggests that the early, high-entropy tokens are sufficient to drive significant alignment with both the teacher’s and the ground-truth distribution.
Left 50% vs. Right 50% For OpenThoughts and Bespoke, training on the first half of each sequence yields markedly lower soft and hard losses than training on the second half. The Left 50% configuration achieves substantially lower loss than the Right 50%, with relative differences consistently negative across different models and loss types. This significant gap can be attributed to (1) early tokens provide stronger learning signal and (2) that the right 50% starts with tokens midway through a reasoning sequence. This abrupt start deprives the model of the necessary context established in the first half, making subsequent token predictions harder and leading to higher losses.
@@
32B-4B Soft -9.20 -6.76 +9.04 +11.65
32B-4B Hard -15.44 -13.06 +8.18 +8.35
8B-4B Soft -14.93 -14.85 +8.99 +10.27
8B-4B Hard -15.68 -13.40 +8.66 +8.76
32B-8B Soft -9.75 -8.14 +9.09 +11.32
32B-8B Hard -16.89 -15.41 +7.43 +7.03
@@
32B-4B Soft -11.73 -14.34 +7.69 +5.86
32B-4B Hard -19.17 -21.49 +6.14 +2.51
8B-4B Soft -15.45 -18.05 +7.97 +7.08
8B-4B Hard -18.76 -21.66 +6.98 +3.63
32B-8B Soft -12.28 -14.93 +7.85 +6.29
32B-8B Hard -20.39 -22.48 +5.61 +2.55
@@
Pair Loss Type Avg (%) Last (%)
32B-4B Soft Loss +12.86 +13.16
32B-4B Hard Loss +9.29 +7.22
8B-4B Soft Loss +14.58 +14.76
8B-4B Hard Loss +10.24 +8.06
32B-8B Soft Loss +12.02 +10.68
32B-8B Hard Loss +6.86 +2.47
Table 9: Relative Difference (%) for Synthetic-1: 50% vs. 100%
Pair Loss Type Avg (%) Last (%)
32B-4B Soft Loss +5.97 +4.56
32B-4B Hard Loss +4.20 +1.72
8B-4B Soft Loss +6.24 +4.95
8B-4B Hard Loss +4.58 +1.88
32B-8B Soft Loss +4.82 +7.16
32B-8B Hard Loss +1.67 +0.85
Table 10: Relative Difference (%) for Nemotron: 50% vs. 100%
Pair Loss Type Avg (%) Last (%)
32B-4B Soft Loss +9.11 +7.68
32B-4B Hard Loss +6.40 +1.95
8B-4B Soft Loss +8.96 +6.66
8B-4B Hard Loss +6.90 +2.14
32B-8B Soft Loss +9.25 +7.60
32B-8B Hard Loss +5.48 +1.03
@@
Figure 19: Loss curves for Synthetic-1: LSP 50% vs LSP 100%
Figure 20: Loss curves for Nemotron: LSP 50% vs LSP 100%
@@
A.4 Ablation Studies
Our ablation objective is to demonstrate that the model performance stems from the method itself rather than specific hyperparameter tuning. We ablate using 8B-4B teacher-student pair and Bespoke dataset among $\lambda\in\{0,0.25,0.5,0.75,1.0\}$. While we adopt $\lambda=0.5$ as the main experimental configuration, Table 12 shows that performance remains consistent across the range $\lambda\in\{0.25,0.5,0.75\}$. The similarity in results among these $\lambda$ values indicates that the choice of 0.5 is not the sole driver of success, but rather that the method performs reliably provided that a balance between hard and soft losses is maintained.
This stability across intermediate $\lambda$ values is accompanied by a clear performance advantage over the pure supervised fine-tuning (SFT) setting ($\lambda=0$). All tested combined configurations ($\lambda\in\{0.25,0.5,0.75\}$) yield higher accuracy than the setting where soft loss is ignored, except for one setting when $\lambda=0.75$ where it is slightly lower than $\lambda=0$ in AIME25 but remains fairly close. Specifically, compared to $\lambda=0$, using our chosen $\lambda=0.5$ results in significant relative improvements: approximately 26.8% (AIME24) and 19.5% (AIME25) for LSP=0.5, and 8.9% (AIME24) and 9.7% (AIME25) for LSP=1.0.
@@ A.4 Ablation Studies
A.5 Miscellaneous
A.5.1 Experimental Results of LSP0.5 vs. LSP1.0
We include the experimental results among five datasets across three teacher-student pairs for LSP0.5 and LSP1.0 on AIME24 and AIME25 in Table 13. For the performance retention rate of LSP0.5 compared to LSP1.0, see Table 4.
Dataset Openthoughts Bespoke Synthetic1 Nemotron SkyT1
AIME24 AIME25 AIME24 AIME25 AIME24 AIME25 AIME24 AIME25 AIME24 AIME25
LSP 0.5 1.0 0.5 1.0 0.5 1.0 0.5 1.0 0.5 1.0 0.5 1.0 0.5 1.0 0.5 1.0 0.5 1.0 0.5 1.0
Pair
32B-4B 22.50 23.39 23.44 24.06 17.66 18.02 19.53 20.94 19.32 24.37 22.08 24.06 25.00 25.89 21.97 24.22 17.71 20.26 18.44 20.57
8B-4B 23.80 24.27 23.91 24.17 16.98 19.79 19.17 21.88 20.89 25.31 22.45 24.37 22.92 25.05 22.19 24.22 17.55 23.33 18.96 21.30
32B-8B 32.19 33.54 27.86 28.54 24.17 26.15 23.33 26.77 26.93 28.28 27.40 28.23 32.08 34.79 26.98 27.92 22.24 27.76 21.35 25.83
Table 13: Experimental results for five datasets in accuracy (%).
A.5.2 Compute Resources
@@
The second research question investigates compute-efficient training via sequence truncation and budget allocation. This involves two sets of experiments. First, we analyze the Sequence-Length Scaling Behavior by varying the Lead-Span Proportion (LSP) across $10$ levels: $p\in\{0.1,0.2,\dots,1.0\}$, representing the fraction of tokens retained from the start of the sequence. This is conducted on all $7$ teacher-student pairs for OpenThoughts and the $3$ largest pairs for Bespoke. Second, we perform a Budget Location Ablation to compare the Left 50% versus Right 50% of tokens. Since the Left 50% configuration is identical to LSP 0.5 (trained and evaluated in the first set), we reuse those results and only perform additional training for the Right 50% configuration. Third, to examine generalizability, we compare LSP 0.5 versus LSP 1.0 on three additional datasets (Synthetic-1, Nemotron, SkyT1) across 3 teacher-student pairs. The total runs ($R_{RQ2}$) are aggregated from the Sequence-Length Scaling ($R_{RQ2}^{\text{Scale}}$), Budget Location Ablation ($R_{RQ2}^{\text{Loc}}$), and Generalizability ($R_{RQ2}^{\text{Gen}}$) components as follows:
@@
$R_{RQ2}^{\text{Gen}}$ $=2\text{ (LSPs)}\times 3\text{ (pairs)}\times 3\text{ (datasets)}=18$
@@
RQ2 124 17,856 8,432 26288
Total 160 23,040 10,880 33,920
A.5.3 Distillation Method Choice